OpenStudy (anonymous):

if logx^8=7/6 what is x please?

OpenStudy (anonymous):

doesnt work

OpenStudy (anonymous):

1.16

OpenStudy (anonymous):

1.15700

OpenStudy (anonymous):

logbasex(8)=7/6

OpenStudy (anonymous):

none of those answer work

OpenStudy (anonymous):

\[\log_{x}8=7/6 \] is it this

OpenStudy (anonymous):

5.94397 is this the ans

OpenStudy (anonymous):

that is correct thanks

OpenStudy (anonymous):

what about logbasex x^3=-4

OpenStudy (anonymous):

logx^8=7/6 8log x=7/6 logx=7/48 x=e^7/48 =1+7/48+(7/48)^2 /2 +.............. =1+.14583333+0.02126736/2 +.......... =1+0.14583333+0.01063368=1.15646701=1.1565(approx)